A= 32 B= 7 A crate with a mass of (12.5+A) kg is lifted by a...

A= 32

B= 7

A crate with a mass of (12.5+A) kg is lifted by a crane. If the crate is accelerating upward at (B/10) m/s2, what is the tension in the cable? Give your answer in newtons (N) and with 3 significant figures.
Answer #1

A = 32 B = M = (12.5 + 32) kg m = 44.5 kg a = 7 = 0.7 m/s² 10 =ma T= mgtma = m (gra) = (44.5 ) ( 9-81 +0.7) = 467.25 Newton